# Predicting Census Survey Response Rates with Parsimonious Additive Models and Structured Interactions

### v4

@@ Abstract
In this paper we consider the problem of predicting survey response rates using a family of flexible and interpretable nonparametric models. The study is motivated by the US Census Bureau’s well-known ROAM application which uses a linear regression model trained on the US Census Planning Database data to identify hard-to-survey areas. A crowdsourcing competition (Erdman and Bates, 2016) organized around ten years ago revealed that machine learning methods based on ensembles of regression trees led to the best performance in predicting survey response rates; however, the corresponding models could not be adopted for the intended application due to their black-box nature. We consider nonparametric additive models with small number of main and pairwise interaction effects using $\ell_{0}$-based penalization. From a methodological viewpoint, we study both computational and statistical aspects of our estimator; and discuss variants that incorporate strong hierarchical interactions. Our algorithms (opensourced on github) extend the computational frontiers of existing algorithms for sparse additive models, to be able to handle datasets relevant for the application we consider. We discuss and interpret findings from our model on the US Census Planning Database. In addition to being useful from an interpretability standpoint, our models lead to predictions that appear to be better than popular black-box machine learning methods based on gradient boosting and feedforward neural networks – suggesting that it is possible to have models that have the best of both worlds: good model accuracy and interpretability.
@@ 1 Introduction
Sample surveys and censuses are primary data sources in social science studies. However, low and often unpredictable response rates in surveys remain a continual source of concern (Erdman and Bates, 2016; Smith, 2014; Tourangeau et al., 2014) – see Figure 1 for an illustration on the American Community Survey. Tourangeau et al. (2014) discuss a multitude of factors that make parts of the population hard-to-survey – such factors are often used to improve sampling strategies, questionnaire designs, recruitment methods, and the language in which the interview is conducted, among others. Erdman and Bates (2016) emphasize the usefulness of having an indicator for hard-to-survey areas to guide targeted surveying (including oversampling), staff recruitment strategies, and targeted follow-ups for non-responders. For major campaigns such as the decennial US Census, this approach can help guide resource allocation for advertisements and building community partnership activities. Eliciting responses from non-self-responding households through follow-up operations can be very costly and time consuming. The Census Bureau estimates that a single percentage increase in the self-response rate amounts to roughly $85$ million dollars saved in personal follow-up costs (US Census Bureau, 2009; Bates and Mulry, 2011). Apart from the cost, the quality of proxy enumerations and imputations is typically significantly lower than that of self-responses (Mule, 2012).
@@
Figure 1: The 2013-2017 American Community Survey self-response rates for all tracts in the continental United States. The North in general, and the Upper Midwest and the Northeast in particular, have higher self-response rates than the rest of the country. Tracts with lower self-response rates are visible in many states – in particular, in the South and in the Mountain region.
Figure 1: The 2013-2017 American Community Survey self-response rates for all tracts in the continental United States. The North in general, and the Upper Midwest and the Northeast in particular, have higher self-response rates than the rest of the country. Tracts with lower self-response rates are visible in many states – in particular, in the South and in the Mountain region.
@@
Both the UK Office for National Statistics and the US Census Bureau have created measures that help quantify the difficulty in gathering data across different geographic areas. A report by Bruce, Robinson and Sanders (2001) from the US Census Bureau introduced a hard-to-count (HTC) score for identifying difficult to enumerate segments of the population. The HTC score, which is based on 12 carefully chosen covariates (for example, housing variables and socio-demographic/economic indicators) has found its use in the planning for the 2010 Census and many other national surveys. The HTC score, however, has some limitations and was later improved to the low-response-score, or LRS (Erdman and Bates, 2016). The LRS plays a key role in the US Census Bureau’s Response Outreach Area Mapper (ROAM) application [^1] https://www.census.gov/library/visualizations/2017/geo/roam.html to identify hard-to-survey areas. The LRS appears to have been partly motivated by the 2012 nationwide competition, organized by the Census Bureau in partnership with Kaggle, on predicting the mail-return rates. This competition aimed to solicit models highly predictive of the decennial census response rates, easily replicable, inherently interpretable, and consistent for use in the field at various levels of geography. Even though the winning models from the competition had good predictive performance, they lacked reproducibility and interpretability. Some of the winning models included covariates that are not publicly available; in particular, they were not chosen from within the US Census Bureau Planning Database. Additionally, the winning models included covariates that were good predictors but not actionable [^2] These include covariates such as the nearest neighboring block group return rates and margins of error for various estimates. Models that incorporate such features are not actionable.. Interpretability suffered due to a multitude of factors. First, the winning models were based on complex ensemble methods (for example, random forests and gradient boosting). Second, these methods employed a large number of covariates (the best model had nearly 340), which hurt the model parsimony. After careful analysis, Erdman and Bates (2016) proposed a linear model based on 25 covariates – LRS is the prediction from this model. While this model suffered in terms of predictive accuracy compared to the black-box machine learning methods, it was highly interpretable and led to useful actionable insights. For example, Erdman and Bates highlight three different block-groups (Columbia Heights, Trinidad, Anacostia) in the District of Columbia that have similar LRS predictions but their characteristics vary, indicating a need for different actions to increase self-response. Columbia Heights has a large Hispanic population, suggesting they could benefit from forms and advertising in Spanish. Anacostia has a stable population (low relocation to the region) and could benefit from community partnerships. On the other hand, Trinidad is characterized as a region in transition with high mobility of the young population. This area could potentially benefit from web-based advertisements for internet-based responses from the tech-savvy younger population.
@@
We aim to predict self-response rates across all Census tracts in the US using operational, socio-economic, and demographic characteristics from the US Census Planning Database. Our goal is to propose interpretable statistical models that deliver strong predictive performance and, thus, can be used to complement the currently used LRS metric. Our hope is that these models lend operational insights into factors influencing survey self-response rates to facilitate the downstream goal of having a cost-effective census with improved coverage.
@@
where $f_{j}$ and $f_{j,k}$ are unknown smooth univariate and bivariate functions, respectively. Drawing inspiration from linear model settings (Bach et al., 2012), we propose new methodology to estimate the unknown functional components via an optimization problem with structural constraints arising from interpretability considerations. To obtain a sparse model with few predictors, that is, with many of the components $\{f_{j}\}$ and $\{f_{j,k}\}$ estimated as exactly zero, we present a novel $\ell_{0}$-regularized approach [^3] For examples of $\ell_{0}$-based approaches and illustrations of their advantages over the $\ell_{1}$-based counterparts in the context of linear regression, see Bertsimas, King and Mazumder (2016); Mazumder and Radchenko (2017); Mazumder, Radchenko and Dedieu (2017); Hazimeh and Mazumder (2020a) and the references therein., penalizing the number of nonzero components in the model. In addition, we explore a refined notion of interpretability in the context of interaction modeling – namely, hierarchical sparsity – inspired by its usage in linear models (for example, Bien, Taylor and Tibshirani, 2013; Hazimeh and Mazumder, 2020b). To our knowledge, we present a novel exploration of computational and statistical perspectives of an $\ell_{0}$ regularized approach for fitting sparse additive models with pairwise interactions.
Applied contributions. Our empirical analysis on the US Census Bureau Planning Database demonstrates the promise of our proposed approach as a potential tool to predict self-response rates in the US Census Bureau application.
@@
Our approach leads to improved prediction of the self-response rates when compared to the linear regression methods discussed in the US Census Bureau report. The improvements are due to both the departure from linearity and the presence of nonlinear interactions. Importantly, the prediction accuracy of our flexible models appears to be at par with (or slightly better than) black-box machine learning methods such as neural networks and gradient boosted decision trees, which topped the nationwide Kaggle competition organized by the Census Bureau.
@@
Our framework results in simple models and allows for automated variable selection. Our models are substantially more compact than many competing benchmarks – we use $8$-$20$ times fewer interaction effects than the corresponding linear models, and about half as many covariates as off-the-shelf machine learning methods (e.g, gradient boosting, feedforward neural networks, or explainable boosting machines) or the sparse linear models with interaction effects. Our models also use a fewer number of covariates (by about 30%) compared to sparse nonparametric additive models (without interactions). We use our models to gather useful operational insights into the factors that influence response rates in surveys across different segments of the population. In particular, our models automatically identify interactions between many of the key factors that have been used in prior publications from the Census Bureau (Bates and Mulry, 2011) to derive meaningful clustering of the population. These interaction effects appear to result in improved predictions of survey response rates and help complement the existing studies by the Census Bureau (Bates and Mulry, 2011; Kulzick et al., 2019) on understanding different segments of the population. In summary, our work appears to address many of the challenges of the US Census application by delivering parsimonious models with good predictive performance.
@@
Methodological contributions. We propose a new family of estimators based on nonparametric additive models with interactions under combinatorial constraints that promote parsimony and interpretability. We establish statistical guarantees for the resulting estimators and present large-scale algorithms to compute these estimators – these algorithms significantly expand the current computational landscape for nonparametric additive models with pairwise interactions. Our approach addresses some of the key computational challenges posed by the large-scale setting of the Census dataset, with $\approx 10^{5}$ nonparametric interaction components and $\approx 10^{5}$ observations. The code implementing our algorithms is available at https://github.com/ShibalIbrahim/Additive-Models-with-Structured-Interactions.
@@
Related Work. There is an impressive body of methodological and theoretical work on using convex $\ell_{1}$-based approaches to fit sparse nonparametric AMs without interactions (see, for example, Meier, van de Geer and Buhlmann, 2009; Ravikumar et al., 2009; Huang, Horowitz and Wei, 2010; Zhao and Liu, 2012; Yuan and Zhou, 2016, and the references therein). Even with main-effects alone, these convex optimization-based methods face computational challenges for the problem-scales we seek to address [^4] Based on our experience, currently available software (for example, R package SAM) encounters numerical difficulties for instances where $n$ is on the order of thousands, and $p$ is on the order of hundreds.. This possibly limits practitioners from realizing the full potential of nonparametric AMs in large-scale settings arising in various practical applications. In terms of statistical properties, the $\ell_{0}$-based estimators can offer improvements over their $\ell_{1}$-counterparts, both on the prediction and the model selection fronts. To this end, we refer the reader to recent work by Hazimeh, Mazumder and Radchenko (2023) demonstrating the advantages of using the $\ell_{0}$-regularized framework for grouped variable selection (which includes nonparametric AMs) without interactions.
@@
In the presence of pairwise interactions, a setting we focus on, the problem of variable selection in nonparametric AMs becomes considerably more challenging. The approaches by Meier, van de Geer and Buhlmann (2009) and Ravikumar et al. (2009) consider additive models with main effects but no pairwise interactions. Lin and Zhang (2006) introduced COSSO, which is a well-known method to fit model (1); however, the method appears to be suitable for low-dimensional settings – for example, the authors consider instances with $n\sim 500$, $p\sim 10$, and $\sim 50$ pairwise interactions [^5] There appears to be no open-source implementation for COSSO..
@@
COSSO penalizes the sum of the Sobolev norms of the functional components in representation (1) and, thus, is capable of producing sparse models. However, the convex COSSO penalty jointly controls sparsity and smoothness, potentially resulting in unwanted shrinkage interfering with component selection. In contrast, our proposed approach uses separate penalties for smoothness and sparsity, and encourages sparsity directly via an $\ell_{0}$-based penalty on the number of components in the model.
@@
More recently, some methods have been proposed that extend the scope of classical nonparametric AMs (Hastie and Tibshirani, 1987). In a series of works Lou et al. (2013); Nori et al. (2019); Yang, Zhang and Sudjianto (2021) explore tree-based and neural-network-based approaches for fitting additive models with sparse interactions. Lou et al. (2013) propose a two-stage approach using shallow tree-like models for the main and interaction effects. In the first stage, they use gradient boosting to fit the main effects – the boosting procedure cycles through all features in a round-robin fashion to fit all main effects. In the second stage, they use a scheme to select a small subset of pairwise interactions–these interaction effects are fitted with shallow tree-like models via gradient boosting. Nori et al. (2019) provide an efficient implementation of the above approach as Explainable Boosting Machines (EBM) in the well-known interpretml toolkit. The EBM procedure is not based on optimizing a penalized estimation framework and, to our knowledge, no statistical guarantees for it are known.
@@
In another line of work, Yang, Zhang and Sudjianto (2021) proposed GAMI-Net, which uses neural networks to fit main and pairwise interaction effects. GAMI-Net uses multi-stage approach: (i) fit all main effects, where each main effect is modeled using a multi-layer perceptron (MLP); (ii) select Top-k interactions based on an interaction detection method (Lou et al., 2013) on the residuals; (iii) fit the Top-k interaction effects simultaneously, where each interaction effect is again modeled as MLP; (iv) fine-tune all selected main effects and interaction effects jointly. GAMI-Net also prunes some components in steps (i) and (iii) based on some ranking measure. Like EBM, the GAMI-Net method does not solve a penalized optimization criterion; no theoretical statistical guarantees for this method are known.
@@
Zschech et al. (2022) independently compared various existing interpretable models and concluded that EBM (Nori et al., 2019) and GAMI-Net (Yang, Zhang and Sudjianto, 2021) appear to be the leading interpretable models. However, both models have some practical limitations. For example, EBM includes all main effects in the model (i.e., there is no feature selection). Neural-based approaches, such as GAMI-Net, are quite computationally expensive. On the Census dataset that we consider, GAMI-Net takes 3 days (using a 8-CPU machine) to compute one modelif we consider 1000 tuning parameters, we are looking at a steep 3000 days of computation cost.
@@
To the best of our knowledge, our algorithmic framework for sparse nonparametric AMs with interactions is novel and is likely to be of independent interest from a computational methodology perspective in applications beyond the one studied herein.
@@
Organization. Section 2 presents the statistical models pursued in this paper. Section 2.3 investigates the statistical properties of our proposed estimators. Section 2.4 discusses how to obtain solutions to the corresponding large-scale discrete optimization problems. Section 3 develops an extension of our method that incorporates strong hierarchy constraints. In Section 4, we present simulation studies comparing our methods with COSSO, EBM and GAMI-Net. Section 5 discusses numerical results for the US Census Bureau application that motivates this study. Additional technical details are provided in the supplementary material.

### v5

@@ Abstract
In this paper, we consider the problem of predicting survey response rates using a family of flexible and interpretable nonparametric models. The study is motivated by the US Census Bureau’s well-known ROAM application, which uses a linear regression model trained on the US Census Planning Database data to identify hard-to-survey areas. A crowdsourcing competition (Erdman and Bates, 2016) organized more than ten years ago revealed that machine learning methods based on ensembles of regression trees led to the best performance in predicting survey response rates; however, the corresponding models could not be adopted for the intended application due to their black-box nature. We consider nonparametric additive models with a small number of main and pairwise interaction effects using $\ell_{0}$-based penalization. From a methodological viewpoint, we study our estimator’s computational and statistical aspects and discuss variants incorporating strong hierarchical interactions. Our algorithms (open-sourced on GitHub) extend the computational frontiers of existing algorithms for sparse additive models to be able to handle datasets relevant to the application we consider. We discuss and interpret findings from our model on the US Census Planning Database. In addition to being useful from an interpretability standpoint, our models lead to predictions comparable to popular black-box machine learning methods based on gradient boosting and feedforward neural networks – suggesting that it is possible to have models that have the best of both worlds: good model accuracy and interpretability.
@@ 1 Introduction
Sample surveys and censuses are primary data sources in social science studies. However, low and often unpredictable response rates in surveys remain a continual source of concern (Erdman and Bates, 2016; Smith, 2014; Tourangeau et al., 2014) – see Figure 1 for an illustration on the American Community Survey. Tourangeau et al. (2014) discuss many factors that make parts of the population hard to survey – such factors are often used to improve sampling strategies, questionnaire designs, recruitment methods, and the language in which the interview is conducted, among others. Erdman and Bates (2016) emphasize the usefulness of having an indicator for hard-to-survey areas to guide targeted surveying (including oversampling), staff recruitment strategies, and targeted nonresponse followup. For major campaigns such as the decennial US Census, this approach can help guide resource allocation for advertisements and building community partnership activities. Eliciting responses from non-self-responding households through follow-up operations can be costly and time-consuming. The Census Bureau estimates that a single percentage increase in the self-response rate amounts to roughly $85$ million dollars saved in personal follow-up costs (US Census Bureau, 2009; Bates and Mulry, 2011). Apart from the cost, the quality of proxy enumerations and imputations is typically substantially lower than that of self-responses (Mule, 2012).
@@
Figure 1: The 2013-2017 American Community Survey self-response rates for all tracts in the continental United States. The North, in general, and the Upper Midwest and the Northeast, in particular, have higher self-response rates than the rest of the country. Tracts with lower self-response rates are visible in many states – in particular, in the South and in the Mountain region.
@@
Both the UK Office for National Statistics and the US Census Bureau have created measures that help quantify the difficulty in gathering data across different geographic areas. A report by Bruce, Robinson and Sanders (2001) from the US Census Bureau introduced a hard-to-count (HTC) score for identifying difficult-to-enumerate segments of the population. The HTC score, based on 12 carefully chosen covariates (for example, housing variables and socio-demographic/economic indicators), has found its use in the planning for the 2010 Census and many other national surveys. The HTC score, however, has some limitations and was later improved to the low-response-score, or LRS (Erdman and Bates, 2016). The LRS plays a key role in the US Census Bureau’s Response Outreach Area Mapper (ROAM) application [^1] https://www.census.gov/library/visualizations/2017/geo/roam.html to identify hard-to-survey areas. The LRS appears to have been partly motivated by the 2012 nationwide competition, organized by the Census Bureau in partnership with Kaggle on predicting the mail-return rates. This competition aimed to solicit models highly predictive of the decennial census response rates, easily replicable, inherently interpretable, and consistent for use in the field at various geographic levels. Even though the winning models from the competition had good predictive performance, they lacked reproducibility and interpretability. Some of the winning models included covariates that are not publicly available; in particular, they were not chosen from within the US Census Bureau Planning Database. Additionally, the winning models included covariates that were good predictors but not actionable [^2] These include covariates such as the nearest neighboring block group return rates and margins of error for various estimates. Models that incorporate such features are not actionable.. Interpretability suffered due to a multitude of factors. First, the winning models were based on complex ensemble methods (e.g., random forests and gradient boosting). Second, these methods employed many covariates (the best model had nearly 340), which hurt the model parsimony. After careful analysis, Erdman and Bates (2016) proposed a linear model based on 25 covariates – LRS is the prediction from this model. While this model suffered loss in predictive accuracy compared to the black-box machine learning methods, it was highly interpretable and led to useful actionable insights. For example, Erdman and Bates highlight three different block groups (Columbia Heights, Trinidad, Anacostia) in the District of Columbia that have similar LRS predictions, but their characteristics vary, indicating a need for different actions to increase self-response. Columbia Heights has a large Hispanic population, suggesting they could benefit from forms and advertising in Spanish. Anacostia has a stable population (with a low relocation rate to the region) and could benefit from community partnerships. On the other hand, Trinidad is characterized as a region in transition with high mobility of the young population. This area could benefit from web-based advertisements for internet-based responses from the tech-savvy younger population.
@@
We aim to predict self-response rates across all Census tracts in the US using operational, socio-economic, and demographic characteristics from the US Census Planning Database. Our goal is to propose interpretable statistical models that deliver strong predictive performance and can potentially complement the currently used LRS metric. Our hope is that these models lend operational insights into factors influencing survey self-response rates to facilitate cost-effectiveness and improved coverage.
@@
where $f_{j}$ and $f_{j,k}$ are unknown smooth univariate and bivariate functions, respectively. Drawing inspiration from linear model settings (Bach et al., 2012), we propose a new methodology to estimate the unknown functional components via an optimization problem with structural constraints arising from interpretability considerations. To obtain a sparse model with few predictors, that is, with many of the components $\{f_{j}\}$ and $\{f_{j,k}\}$ estimated as exactly zero, we present a novel $\ell_{0}$-regularized approach [^3] For examples of $\ell_{0}$-based approaches and illustrations of their advantages over the $\ell_{1}$-based counterparts in the context of linear regression, see Bertsimas, King and Mazumder (2016); Mazumder and Radchenko (2017); Mazumder, Radchenko and Dedieu (2023); Hazimeh and Mazumder (2020a); Hazimeh, Mazumder and Saab (2022a) and the references therein., penalizing the number of nonzero components in the model. In addition, we explore a refined notion of interpretability in the context of interaction modeling – namely, hierarchical sparsity – inspired by its usage in linear models (for example, Bien, Taylor and Tibshirani, 2013; Hazimeh and Mazumder, 2020b). To our knowledge, we present a novel exploration of computational and statistical perspectives of an $\ell_{0}$ regularized approach for fitting sparse additive models with pairwise interactions.
@@
Applied contributions. Our empirical analysis of the US Census Bureau Planning Database demonstrates the promise of our proposed approach. Our estimator improves the prediction of the self-response rates when compared to the linear regression methods discussed in the US Census Bureau report. The improvements are due to both the departure from linearity and the presence of nonlinear interactions. Importantly, the prediction accuracy of our flexible models appears to be at par with (or slightly better than) black-box machine learning methods such as neural networks and gradient-boosted decision trees, which topped the nationwide Kaggle competition organized by the Census Bureau.
@@
Our framework results in simple models and allows for automated variable selection. Our models are substantially more compact than many competing benchmarks – we use $8$-$20$ times fewer interaction effects than the corresponding linear models and about half as many covariates as off-the-shelf machine learning methods (e.g., gradient boosting, feedforward neural networks, or explainable boosting machines) or the sparse linear models with interaction effects. Our models also use fewer covariates (by about 30%) compared to sparse nonparametric additive models (without interactions)—see Section 5 for details. We use our models to gather useful operational insights into the factors that influence response rates in surveys across different segments of the population. In particular, our models automatically identify interactions between many of the key factors used in prior publications from the Census Bureau (Bates and Mulry, 2011) to derive meaningful population clustering. These interaction effects appear to result in improved predictions of survey response rates and complement earlier findings reported in Bates and Mulry (2011); Kulzick et al. (2019) on understanding different population segments.
@@
Methodological contributions. We propose a new family of estimators based on nonparametric additive models with sparse interactions under combinatorial constraints. We establish statistical guarantees for the resulting estimators and present large-scale algorithms for these estimators extending the current computational landscape for sparse nonparametric additive models with pairwise interactions. Our approach addresses some of the key computational challenges posed by the large-scale setting of the Census dataset, with approx. $10^{5}$ nonparametric interaction components and approx. $10^{5}$ observations. The code implementing our algorithms is available at:
Code: https://github.com/mazumder-lab/elaan.
@@
Related Work. There is an impressive body of methodological and theoretical work on using convex $\ell_{1}$-based approaches to fit sparse nonparametric AMs without interactions (see, for example, Meier, van de Geer and Buhlmann, 2009; Ravikumar et al., 2009; Huang, Horowitz and Wei, 2010; Zhao and Liu, 2012; Yuan and Zhou, 2016, and the references therein). Even with main-effects alone, these convex optimization-based methods face computational challenges for the problem-scales we seek to address [^4] Based on our experience, currently available software (for example, R package SAM) encounters numerical difficulties for instances where $n$ is on the order of thousands, and $p$ is on the order of hundreds.. This possibly limits practitioners from realizing the full potential of nonparametric AMs in large-scale settings arising in various practical applications. Regarding statistical properties, the $\ell_{0}$-based estimators can offer improvements over their $\ell_{1}$-counterparts, both on the prediction and the model selection fronts. To this end, we refer the reader to recent work by Hazimeh, Mazumder and Radchenko (2023) demonstrating the advantages of using the $\ell_{0}$-regularized framework for grouped variable selection (which includes nonparametric AMs) without interactions.
@@
In the presence of pairwise interactions, a setting we focus on, the problem of variable selection in nonparametric AMs becomes considerably more challenging. The approaches by Meier, van de Geer and Buhlmann (2009) and Ravikumar et al. (2009) consider additive models with main effects but no pairwise interactions. Lin and Zhang (2006) introduced COSSO, which is a well-known method to fit model (1); however, the method appears to be suitable for low-dimensional settings – for example, the authors consider instances with $n\sim 500$, $p\sim 10$, and approximately $50$ pairwise interactions [^5] There appears to be no open-source implementation for COSSO..
@@
COSSO penalizes the sum of the Sobolev norms of the functional components in representation (1) and, thus, can produce sparse models. However, the convex COSSO penalty jointly controls sparsity and smoothness, potentially resulting in unwanted shrinkage interfering with component selection. In contrast, our proposed approach uses separate penalties for smoothness and sparsity, and encourages sparsity directly via an $\ell_{0}$-based penalty on the number of components in the model.
@@
More recently, there have been approaches that extend the scope of classical nonparametric AMs (Hastie and Tibshirani, 1987). In a series of works Lou et al. (2013); Nori et al. (2019); Yang, Zhang and Sudjianto (2021) explore tree-based and neural-network-based approaches for fitting additive models with sparse interactions. Lou et al. (2013) propose a two-stage approach using shallow tree-like models for the main and interaction effects. In the first stage, they use gradient boosting to fit the main effects – the boosting procedure cycles through all features in a round-robin fashion to fit all main effects. In the second stage, they use a scheme to select a small subset of pairwise interactions–these interaction effects are fitted with shallow tree-like models via gradient boosting. Nori et al. (2019) provide an efficient implementation of the above approach as Explainable Boosting Machines (EBM) in the well-known interpretml toolkit. The EBM procedure is not based on optimizing a penalized estimation framework and, to our knowledge, no statistical guarantees for it are known.
@@
In another line of work, Yang, Zhang and Sudjianto (2021) proposed GAMI-Net, which uses neural networks to fit main and pairwise interaction effects. GAMI-Net uses a multi-stage approach: (i) fit all main effects, where each main effect is modeled using a multi-layer perceptron (MLP); (ii) select Top-k interactions based on an interaction detection method (Lou et al., 2013) on the residuals; (iii) fit the Top-k interaction effects simultaneously, where each interaction effect is again modeled as MLP; (iv) fine-tune all selected main effects and interaction effects together. GAMI-Net also prunes some components in steps (i) and (iii) based on some ranking measure. GAMI-Net uses screening methods prior to training to select a collection of main and interaction effects and, as such, does not jointly optimize the sparsity pattern while training.
@@
Zschech et al. (2022) compared various additive models and observed that EBM (Nori et al., 2019) and GAMI-Net (Yang, Zhang and Sudjianto, 2021) led to good performance. However, both models have some practical limitations. For example, EBM includes all main effects in the model (i.e., there is no feature selection). Neural-based approaches, such as GAMI-Net, are quite computationally expensive. On the Census dataset that we consider, GAMI-Net takes 12 hours (using an 8-CPU machine) to compute one model; thus, using 1000 tuning parameters leads to a steep 500-day computational cost  [^6] This time is for the architecture that uses a 2-layered NN with 10 neurons for each main/interaction effect..
@@
To the best of our knowledge, our algorithmic framework for sparse nonparametric AMs with interactions is novel and is likely to be of independent interest from a computational methodology perspective.
@@
Organization. Section 2 presents the statistical models pursued in this paper. Section 2.3 investigates the statistical properties of our proposed estimators. Section 2.4 discusses how to obtain solutions to the corresponding large-scale discrete optimization problems. Section 3 develops an extension of our method that incorporates strong hierarchy constraints. In Section 4, we present simulation studies comparing our methods with COSSO, EBM and GAMI-Net. Section 5 discusses numerical results for the US Census Bureau application that motivates this study. Additional technical details are provided in the supplementary material (Ibrahim et al., 2024).